# TMI! Finetuned Models Leak Private Information from their Pretraining Data

### v2

@@ 4.2 Adapting an Existing Attack
As a first attempt to create an effective membership-inference attack on finetuned machine learning models, we can consider an adaptation of the likelihood ratio attack (LiRA) proposed by Carlini et al. [13]. In this attack (Algorithm 2), the adversary observes the target model’s prediction confidence on a challenge point with respect to the true label of the challenge point. Because the model’s confidence with respect to a given label is approximately normally distributed, Carlini et al. perform a likelihood ratio test to infer the challenge point’s membership status, using a set of shadow models to parameterize the IN and OUT distributions.
In our setting, these shadow models are first trained on datasets drawn from $\mathbb{D}_{PT}$, then finetuned on a dataset drawn from $\mathbb{D}_{FT}$ (Algorithm 1). Because the ground truth label of the challenge point drawn from $\mathbb{D}_{PT}$ is not necessarily in the range of our finetuned target model we cannot perform the likelihood ratio test with respect to the observed confidence on the point’s true label. Instead, we can adapt the attack to use the label predicted by the target model with the highest confidence, $\hat{y}$. To do this, we store the entire prediction vector for each query to our shadow models, and only use the scaled model confidences at index $\hat{y}$, denoted $f(x)_{\hat{y}}$, of the prediction vectors. We follow the lead of Carlini et al. [13] and query each shadow and target model on $M$ random augmentations of the challenge point and fit $M$-dimensional multivariate normal distributions to the scaled model confidences we aggregate to improve attack success.
@@
We adapt the MI attack shown in [13] by using the label which the target model predicted most confidently instead of the ground truth label.
@@ 4.3 Issues with Adapting LiRA
While this adaptation of LiRA is somewhat effective at inferring membership (Figures 5 and 6), it only captures how the pretraining dataset influences model’s predictions with respect to a single label in the downstream dataset. Because the purpose of pretraining is to extract and learn general features that can be used in several downstream tasks, one would expect that the weights of a pretrained model have some impact on all of a finetuned model’s prediction confidences. For example, Figure 4 shows that the presence of a specific image labeled as "dugong" in the training set makes finetuned models, which cannot themselves predict the label "dugong", more confident on their downstream prediction of "elephant" and "platypus". Meanwhile, the presence of this image in the training dataset has little to no impact on the downstream label "scissors".
@@
Figure 4: Scaled Model Confidences of Shadow Models Finetuned on Caltech 101 at Multiple Labels when Queried on a Sample from the "Dugong" Class in Tiny Imagenet
@@ 4.4 Our TMI Attack
In our implementation of TMI for computer vision models, we train a metaclassifier per challenge point. Because we use a relatively small number of shadow models (64 IN and 64 OUT in total), we leverage random augmentations to construct a larger metaclassifier dataset. Each time we query the target model or our local shadow models, we query $M$ times with different random augmentations of the challenge point, including random horizontal flips and random crops with padding. This yields $M\times 2\times 64$ prediction vectors for each challenge point. In total, our metaclassifiers are trained on 1024 labeled prediction vectors, 512 labeled 00 to denote "non-member" or OUT and 512 labeled $1$ to denote "member" or IN.
@@ 5 TMI Evaluation
We evaluate the performance of our TMI attack on image models with two pretraining tasks and four downstream tasks and public, pretrained language models with one pretraining task and two downstream tasks. We evaluate the success of our attack as a function of the number of updated parameters, and we choose downstream tasks with differing similarity to the pretraining task to show how attack success depends on the relevance of the pretraining task to the downstream task. Additionally, we observe the success of our attack when differential privacy [23] is used in the finetuning process, which is an increasingly popular technique to maintain utility while preserving the privacy of individuals in the dataset of downstream task [64, 12, 6, 31, 29, 42].
@@ 5.1.1 Datasets
We pretrain our small vision models on CIFAR-100 [39] and finetune them on a coarse-labeled version of CIFAR-100, CIFAR-10 [39], and Oxford-IIIT Pet [49]. Our larger vision models are pretrained on Tiny ImageNet [56] and finetuned on Caltech 101 [41]. For our language tasks, we use publicly available pretrained WikiText-103 [46] models and finetune them on DBpedia [67] and Yahoo Answers [68] topic classification datasets. A detailed description of the datasets used in our evaluation can be found in Appendix C.1.
@@ 5.1.2 Models
For our vision tasks, we use the ResNet-34 [32] and Wide ResNet-101 [65] architectures. The ResNet architecture has been widely used in various computer vision applications due to its superior performance and efficiency. ResNet is a convolutional neural network architecture that uses residual blocks, allowing it to effectively handle the complex features of images and perform well on large-scale datasets.
For our language tasks, we use the Transformer-XL [20] model architecture. In particular, we use the pretrained Transformer-XL model from Hugging Face, which is trained on WikiText-103 [46], as our initialization for the downstream tasks. We finetune our pretrained language model architectures on the DBpedia ontology classification and Yahoo Answers topic classification datasets.
@@ 5.2 Metrics
To evaluate the performance of TMI, we use a set of metrics that are commonly used in the literature. The first metric is balanced attack accuracy, which measures the percentage of samples for which our attack correctly identifies membership status. Although balanced accuracy is a common metric used to evaluate MI attacks [53, 18, 51, 63], prior work [13] argues that it is not sufficient by itself to measure the performance of MI attacks as privacy is not an average case metric [55]. Therefore, we also evaluate our attack using the receiver operating characteristic (ROC) curve.
@@ 5.3 Experimental Results
During training, we designate 1000 and 2000 samples to be challenge points for CIFAR-100 and Tiny ImageNet, respectively, and we run our attack for each of these challenge points on 128 different target models. We compare our results to performing LiRA [13] directly on the pretrained model (i.e., the adversary has access to the model before it was finetuned) to provide an upper bound on our attack’s performance.
@@ 5.3.2 Updating Model Parameters
Prior work [37] has shown that samples used earlier in training are more robust to privacy attacks. Our theoretical results in Section 4.1 substantiate this work and help provide an explanation for the decrease in attack success. In our results, $\alpha$ corresponds to the fraction of training epochs spent on finetuning, but our analysis lacks a critical parameter from our experiments: the number or fraction of tunable parameters in the published statistic. Our analysis considers a vector (namely, the empirical mean) where all of the parameters are being updated, thus providing a worst-case situation for the adversary. In the feature extraction setting, the information learned by the model during pretraining is essentially frozen. Unlike feature extraction, we are updating the model’s parameters with information about the downstream samples when finetuning.
@@ 5.3.3 Finetuning Pretrained Language Models
To answer this research question, we evaluate the success of our TMI attack in the natural language domain. In particular, we focus on publicly available pretrained large language models (LLMs), or foundation models [9], which we finetune on two text classification tasks.
Due to computational limitations, we do not train LLMs from scratch. As an alternative, we evaluate our attack on a widely used pretrained foundation model, Transformer-XL [20], along with its corresponding tokenizer, which are hosted by Hugging Face [59]. Only a limited number of organizations with sufficient computational resources possess the capability to train foundation models, which are typically fine-tuned on specific tasks by smaller organizations [2, 3, 4]. Through our evaluation of TMI on finetuned foundation models, we will additionally answer the following research question:
@@
We chose this foundation model in particular because it uses known training, validation, and testing splits from the WikiText-103 [46] dataset, providing us with the exact partitions necessary to evaluate TMI without having to train our own LLMs. Additionally, although modest in comparison to contemporary foundation models, the Transformer-XL architecture contains 283 million trainable parameters. This makes it a powerful and expressive language model that may be prone to memorizing individual data points.
We finetune Transformer-XL on DBpedia [67], modifying the pretrained tokenizer to use a max length of 450, including both truncation and padding. Using a training set of 10,000 randomly sampled datapoints from DBpedia, we finetune the last third of the parameters in our Tranformer-XL models for 1 epoch. We use the AdamW [44] optimizer with a learning rate of $10^{-5}$ and weight decay with $\lambda=10^{-5}$. With these hyperparameters, we are able to achieve test accuracies of $97\%$ and $60\%$ on the 14 classes of DBpedia and 10 classes of Yahoo Answers, respectively.
@@ 5.3.4 Transfer Learning with Differential Privacy
We also discuss the performance of our attack on target models that were finetuned with differential privacy. Because prior work on transfer learning with differential privacy considers strategies where an especially small percentage of parameters are trained on the downstream task [64, 12, 6, 48], we freeze the pretrained model’s weights and train only the final layer on the downstream task. In our experiments, we perform feature extraction to finetune our pretrained CIFAR-100 models on Coarse CIFAR-100 and CIFAR-10. We train the final classification layer using DP-SGD [6] with target privacy parameters $\varepsilon=\{0.5,1\}$ and $\delta=10^{-5}$. As these are strict privacy parameters, we set the clipping norm equal to $5$ to achieve reasonable utility on the downstream tasks.
@@
Prior work [13] has shown that state-of-the-art MI attacks, which directly query the pretrained model, completely fail when the target models are trained with a small amount of additive noise. For example, when training target models using DP-SGD with a clipping norm equal to $5$ and privacy parameter $\varepsilon=8$, LiRA has an AUC of 0.5. Through this evaluation, we reinforce the fact that transferring pretrained models to downstream tasks with differential privacy does not provide a privacy guarantee for the pretraining data.
@@
While it may seem expected that finetuning on a disjoint dataset with DP-SGD provides no privacy guarantee for individuals in the pretraining set, the authors of [57] pose the following question: What privacy guarantee should an individual expect if their data was present in both pretraining and finetuning? This scenario is not unlikely, as large models are trained on terabytes of data scraped from the Web [11, 60]. Because manually inspecting these datasets is infeasible, it is likely that an individual’s datapoint which was included in private finetuning is also present in non-private pretraining. Thus, their data does not enjoy the $(\varepsilon,\delta)-DP$ guarantee promised by finetuning, as the corresponding pretraining gradients are unbounded in magnitude and exact in direction. Misusing DP-SGD in this manner can leave these individuals at risk of privacy attacks.
To support our claim that finetuning with DP-SGD is blatantly non-private when $D_{PT}\cap D_{FT}\neq\emptyset$, we run experiments on the CIFAR-100 dataset. Similar to our prior experiments, we finetune the final layer of ResNet-34 shadow models on the Coarse CIFAR-100 dataset using DP-SGD [6] with target privacy parameters $\varepsilon\in\{0.5,1\}$ and $\delta=10^{-5}$ and clipping norm equal to 5. The only difference in this experiment is the fact that the finetuning set contains some ($\sim$1000) individuals who were also present in the pretraining task. We run our TMI attack on these individuals and report the results in Figure 13 and Table 7. When we finetune with DP-SGD and the challenge points are included in the finetuning set, we see true positive rates that are comparable to our experiments on models finetuned using feature extraction (Table 1). At a fixed FPR of 0.1% and target privacy guarantees of $\varepsilon=0.5$ and $\varepsilon=1$, TMI achieves a TPR 15.6$\times$ higher than the upper bound (end-to-end) training with DP-SGD should provide.
@@ 6 Discussion and Conclusion
We introduce the first threat model that uses query access to a finetuned model to mount a privacy attack on pretraining data. It remains an open question as to whether other privacy attacks, such as property inference, attribute inference, and training data extraction attacks can also see success in this transfer learning setting. Given that MI attacks are used as practical tools to measure or audit the privacy of machine learning models [54, 35, 62], future work should consider efficiency and simplicity when designing new privacy attacks in the transfer learning setting.
@@
Our evaluation shows that the pretraining dataset of machine learning models finetuned with differential privacy are still susceptible to privacy leakage. This supports the argument made in [57] that "privacy-preserving" models derived from large, pretrained models don’t necessarily provide the privacy guarantees that consumers of services backed by these finetuned models would expect. Prior works that utilize public data to improve the utility of differentially private machine learning models have made strides towards making differential privacy practical for several deep learning tasks [48, 64, 42, 12, 31, 29, 30], but they do not address privacy risks external to model training itself.
Using TMI as a measurement of privacy leakage in this setting, we reinforce the fact that maintaining privacy depends on taking a holistic approach to the way that training data is handled. As stated in [57], privacy is not binary (i.e. not all data is either strictly "private" or "public") and privacy in machine learning is not only dependent on the model’s training procedure. To grapple with privacy risk in this increasingly popular transfer learning setting, researchers and practitioners should explore new ways to sanitize sensitive information from training datasets of machine learning models, create ways to collect potentially sensitive Web data with informed consent from individuals, and work towards end-to-end privacy-preserving machine learning with high utility and privacy guarantees.

### v3

@@ 4.2 Adapting an Existing Attack
As a first attempt to create an effective membership-inference attack on finetuned machine learning models, we can consider an adaptation of the likelihood ratio attack (LiRA) proposed by Carlini et al. [carlini-2022-lira]. In this attack (Algorithm 2), the adversary observes the target model’s prediction confidence on a challenge point with respect to the true label of the challenge point. Because the model’s confidence with respect to a given label is approximately normally distributed, Carlini et al. perform a likelihood ratio test to infer the challenge point’s membership status, using a set of shadow models to parameterize the IN and OUT distributions.
In our setting, these shadow models are first trained on datasets drawn from $\mathbb{D}_{PT}$, then finetuned on a dataset drawn from $\mathbb{D}_{FT}$ (Algorithm 1). Because the ground truth label of the challenge point drawn from $\mathbb{D}_{PT}$ is not necessarily in the range of our finetuned target model we cannot perform the likelihood ratio test with respect to the observed confidence on the point’s true label. Instead, we can adapt the attack to use the label predicted by the target model with the highest confidence, $\hat{y}$. To do this, we store the entire prediction vector for each query to our shadow models, and only use the scaled model confidences at index $\hat{y}$, denoted $f(x)_{\hat{y}}$, of the prediction vectors. We follow the lead of Carlini et al. [carlini-2022-lira] and query each shadow and target model on $M$ random augmentations of the challenge point and fit $M$-dimensional multivariate normal distributions to the scaled model confidences we aggregate to improve attack success.
@@
We adapt the MI attack shown in [carlini-2022-lira] by using the label which the target model predicted most confidently instead of the ground truth label.
@@ 4.3 Issues with Adapting LiRA
While this adaptation of LiRA is somewhat effective at inferring membership (Figures 5 and 6), it only captures how the pretraining dataset influences model’s predictions with respect to a single label in the downstream dataset. Because the purpose of pretraining is to extract and learn general features that can be used in several downstream tasks, one would expect that the weights of a pretrained model have some impact on all of a finetuned model’s prediction confidences. For example, Figure 4 shows that the presence of a specific image labeled as ”dugong” in the training set makes finetuned models, which cannot themselves predict the label ”dugong”, more confident on their downstream prediction of ”elephant” and ”platypus”. Meanwhile, the presence of this image in the training dataset has little to no impact on the downstream label ”scissors”.
@@
Figure 4: Scaled Model Confidences of Shadow Models Finetuned on Caltech 101 at Multiple Labels when Queried on a Sample from the ”Dugong” Class in Tiny Imagenet
@@ 4.4 Our TMI Attack
In our implementation of TMI for computer vision models, we train a metaclassifier per challenge point. Because we use a relatively small number of shadow models (64 IN and 64 OUT in total), we leverage random augmentations to construct a larger metaclassifier dataset. Each time we query the target model or our local shadow models, we query $M$ times with different random augmentations of the challenge point, including random horizontal flips and random crops with padding. This yields $M\times 2\times 64$ prediction vectors for each challenge point. In total, our metaclassifiers are trained on 1024 labeled prediction vectors, 512 labeled 00 to denote ”non-member” or OUT and 512 labeled $1$ to denote ”member” or IN.
@@ 5 TMI Evaluation
We evaluate the performance of our TMI attack on image models with two pretraining tasks and four downstream tasks and public, pretrained language models with one pretraining task and two downstream tasks. We evaluate the success of our attack as a function of the number of updated parameters, and we choose downstream tasks with differing similarity to the pretraining task to show how attack success depends on the relevance of the pretraining task to the downstream task. Additionally, we observe the success of our attack when differential privacy [dwork-2006-dmns] is used in the finetuning process, which is an increasingly popular technique to maintain utility while preserving the privacy of individuals in the dataset of downstream task [yu-2022-dpllm, bu-2023-dpbiasft, abadi-2016-dpsgd, he-2022-groupwisedp, ganesh-2023-pretraining, li-2022-llmstrongdp].
@@ 5.1.1 Datasets
We pretrain our small vision models on CIFAR-100 [CIFAR] and finetune them on a coarse-labeled version of CIFAR-100, CIFAR-10 [CIFAR], and Oxford-IIIT Pet [pets]. Our larger vision models are pretrained on Tiny ImageNet [tiny-imagenet] and finetuned on Caltech 101 [caltech-101]. For our language tasks, we use publicly available pretrained WikiText-103 [wikitext103] models and finetune them on DBpedia [dbpedia] and Yahoo Answers [2015-yahoo] topic classification datasets. A detailed description of the datasets used in our evaluation can be found in Appendix C.1.
@@ 5.1.2 Models
For our vision tasks, we use the ResNet-34 [ResNet] and Wide ResNet-101 [wide-resnet] architectures. The ResNet architecture has been widely used in various computer vision applications due to its superior performance and efficiency. ResNet is a convolutional neural network architecture that uses residual blocks, allowing it to effectively handle the complex features of images and perform well on large-scale datasets.
For our language tasks, we use the Transformer-XL [transformerxl] model architecture. In particular, we use the pretrained Transformer-XL model from Hugging Face, which is trained on WikiText-103 [wikitext103], as our initialization for the downstream tasks. We finetune our pretrained language model architectures on the DBpedia ontology classification and Yahoo Answers topic classification datasets.
@@ 5.2 Metrics
To evaluate the performance of TMI, we use a set of metrics that are commonly used in the literature. The first metric is balanced attack accuracy, which measures the percentage of samples for which our attack correctly identifies membership status. Although balanced accuracy is a common metric used to evaluate MI attacks [shokri-2016-mi, choquettechoo-2021-labelonly, sablayrolles-2019-whitebox, 2018-yeom-mioverfitting], prior work [carlini-2022-lira] argues that it is not sufficient by itself to measure the performance of MI attacks as privacy is not an average case metric [steinke-2020-averageblog]. Therefore, we also evaluate our attack using the receiver operating characteristic (ROC) curve.
@@ 5.3 Experimental Results
During training, we designate 1000 and 2000 samples to be challenge points for CIFAR-100 and Tiny ImageNet, respectively, and we run our attack for each of these challenge points on 128 different target models. We compare our results to performing LiRA [carlini-2022-lira] directly on the pretrained model (i.e., the adversary has access to the model before it was finetuned) to provide an upper bound on our attack’s performance.
@@ 5.3.2 Updating Model Parameters
Prior work [jagielski-2023-forgetting] has shown that samples used earlier in training are more robust to privacy attacks. Our theoretical results in Section 4.1 substantiate this work and help provide an explanation for the decrease in attack success. In our results, $\alpha$ corresponds to the fraction of training epochs spent on finetuning, but our analysis lacks a critical parameter from our experiments: the number or fraction of tunable parameters in the published statistic. Our analysis considers a vector (namely, the empirical mean) where all of the parameters are being updated, thus providing a worst-case situation for the adversary. In the feature extraction setting, the information learned by the model during pretraining is essentially frozen. Unlike feature extraction, we are updating the model’s parameters with information about the downstream samples when finetuning.
@@ 5.3.3 Finetuning Pretrained Language Models
To answer this research question, we evaluate the success of our TMI attack in the natural language domain. In particular, we focus on publicly available pretrained large language models (LLMs), or foundation models [foundationmodels], which we finetune on two text classification tasks.
Due to computational limitations, we do not train LLMs from scratch. As an alternative, we evaluate our attack on a widely used pretrained foundation model, Transformer-XL [transformerxl], along with its corresponding tokenizer, which are hosted by Hugging Face [huggingface]. Only a limited number of organizations with sufficient computational resources possess the capability to train foundation models, which are typically fine-tuned on specific tasks by smaller organizations [openai-api, google-api, microsoft-api]. Through our evaluation of TMI on finetuned foundation models, we will additionally answer the following research question:
@@
We chose this foundation model in particular because it uses known training, validation, and testing splits from the WikiText-103 [wikitext103] dataset, providing us with the exact partitions necessary to evaluate TMI without having to train our own LLMs. Additionally, although modest in comparison to contemporary foundation models, the Transformer-XL architecture contains 283 million trainable parameters. This makes it a powerful and expressive language model that may be prone to memorizing individual data points.
We finetune Transformer-XL on DBpedia [dbpedia], modifying the pretrained tokenizer to use a max length of 450, including both truncation and padding. Using a training set of 10,000 randomly sampled datapoints from DBpedia, we finetune the last third of the parameters in our Tranformer-XL models for 1 epoch. We use the AdamW [loshchilov-2017-adamw] optimizer with a learning rate of $10^{-5}$ and weight decay with $\lambda=10^{-5}$. With these hyperparameters, we are able to achieve test accuracies of $97\%$ and $60\%$ on the 14 classes of DBpedia and 10 classes of Yahoo Answers, respectively.
@@ 5.3.4 Transfer Learning with Differential Privacy
We also discuss the performance of our attack on target models that were finetuned with differential privacy. Because prior work on transfer learning with differential privacy considers strategies where an especially small percentage of parameters are trained on the downstream task [yu-2022-dpllm, bu-2023-dpbiasft, abadi-2016-dpsgd, papernot-2020-making], we freeze the pretrained model’s weights and train only the final layer on the downstream task. In our experiments, we perform feature extraction to finetune our pretrained CIFAR-100 models on Coarse CIFAR-100 and CIFAR-10. We train the final classification layer using DP-SGD [abadi-2016-dpsgd] with target privacy parameters $\varepsilon=\{0.5,1\}$ and $\delta=10^{-5}$. As these are strict privacy parameters, we set the clipping norm equal to $5$ to achieve reasonable utility on the downstream tasks.
@@
Prior work [carlini-2022-lira] has shown that state-of-the-art MI attacks, which directly query the pretrained model, completely fail when the target models are trained with a small amount of additive noise. For example, when training target models using DP-SGD with a clipping norm equal to $5$ and privacy parameter $\varepsilon=8$, LiRA has an AUC of 0.5. Through this evaluation, we reinforce the fact that transferring pretrained models to downstream tasks with differential privacy does not provide a privacy guarantee for the pretraining data.
@@
While it may seem expected that finetuning on a disjoint dataset with DP-SGD provides no privacy guarantee for individuals in the pretraining set, the authors of [2022-tramer-position] pose the following question: What privacy guarantee should an individual expect if their data was present in both pretraining and finetuning? This scenario is not unlikely, as large models are trained on terabytes of data scraped from the Web [2020-gpt3, JFT300]. Because manually inspecting these datasets is infeasible, it is likely that an individual’s datapoint which was included in private finetuning is also present in non-private pretraining. Thus, their data does not enjoy the $(\varepsilon,\delta)-DP$ guarantee promised by finetuning, as the corresponding pretraining gradients are unbounded in magnitude and exact in direction. Misusing DP-SGD in this manner can leave these individuals at risk of privacy attacks.
To support our claim that finetuning with DP-SGD is blatantly non-private when $D_{PT}\cap D_{FT}\neq\emptyset$, we run experiments on the CIFAR-100 dataset. Similar to our prior experiments, we finetune the final layer of ResNet-34 shadow models on the Coarse CIFAR-100 dataset using DP-SGD [abadi-2016-dpsgd] with target privacy parameters $\varepsilon\in\{0.5,1\}$ and $\delta=10^{-5}$ and clipping norm equal to 5. The only difference in this experiment is the fact that the finetuning set contains some ($\sim$1000) individuals who were also present in the pretraining task. We run our TMI attack on these individuals and report the results in Figure 13 and Table 7. When we finetune with DP-SGD and the challenge points are included in the finetuning set, we see true positive rates that are comparable to our experiments on models finetuned using feature extraction (Table 1). At a fixed FPR of 0.1% and target privacy guarantees of $\varepsilon=0.5$ and $\varepsilon=1$, TMI achieves a TPR 15.6$\times$ higher than the upper bound (end-to-end) training with DP-SGD should provide.
@@ 6 Discussion and Conclusion
We introduce the first threat model that uses query access to a finetuned model to mount a privacy attack on pretraining data. It remains an open question as to whether other privacy attacks, such as property inference, attribute inference, and training data extraction attacks can also see success in this transfer learning setting. Given that MI attacks are used as practical tools to measure or audit the privacy of machine learning models [song-2018-auditingmi, tensorflow-2020-auditing, ye-2022-enhanced_mi], future work should consider efficiency and simplicity when designing new privacy attacks in the transfer learning setting.
@@
Our evaluation shows that the pretraining dataset of machine learning models finetuned with differential privacy are still susceptible to privacy leakage. This supports the argument made in [2022-tramer-position] that ”privacy-preserving” models derived from large, pretrained models don’t necessarily provide the privacy guarantees that consumers of services backed by these finetuned models would expect. Prior works that utilize public data to improve the utility of differentially private machine learning models have made strides towards making differential privacy practical for several deep learning tasks [papernot-2020-making, yu-2022-dpllm, li-2022-llmstrongdp, bu-2023-dpbiasft, he-2022-groupwisedp, ganesh-2023-pretraining, golatkar-2022-mixeddp], but they do not address privacy risks external to model training itself.
Using TMI as a measurement of privacy leakage in this setting, we reinforce the fact that maintaining privacy depends on taking a holistic approach to the way that training data is handled. As stated in [2022-tramer-position], privacy is not binary (i.e. not all data is either strictly ”private” or ”public”) and privacy in machine learning is not only dependent on the model’s training procedure. To grapple with privacy risk in this increasingly popular transfer learning setting, researchers and practitioners should explore new ways to sanitize sensitive information from training datasets of machine learning models, create ways to collect potentially sensitive Web data with informed consent from individuals, and work towards end-to-end privacy-preserving machine learning with high utility and privacy guarantees.
[…]